Assume that when adults with smartphones are randomly​ selected, 52% use them in meetings or classes. If 30 adults smartphones users are randomly selected, find the probability that exactly 24 of them use their smartphone in meetings or classes.

The probability is?
Answer #1

Here we have : P=0.52 n=30 Binomial distribution ! P(x+x)=nex p (1-P)1-3 p(8=24) =30(24(0.52)2 (1-0-52/30-24 = 10.00111